What is the pH of a 0.25 M solution of anilinium nitrate, C6H5NH3NO3? Kb for aniline is 4.3 × 10-10.
Answer #1

Answer :- Kb = 4-3x1610 Coloring - o & to Chrik Initial 0.25m 0 O change -X + X - X At eam 0.25-X X Х ka ko =kw -14 = I xio k